a mixture of two isotopes 107 ag with a mass of 106.9041 and 109 ag with a mass of 108.9047 the average atomic wegith of silver is 107.868 what are the relative abudances expressed as precents of these two silver isotopes ?


1
Expert's answer
2021-04-29T08:23:25-0400

106.905.0.5184+0.4816=107.87

107.87-55.419552=0.4816

52.450448/0.4816=108.909

=108.909